OpenStudy (anonymous):

Adam and Gloria traveled in a sailboat for 4 hours with a 6 mph current in the water. Against the same current, it took them 6 hours to return home. How fast was the boat traveling in still water?

OpenStudy (anonymous):

@johnweldon1993 can you help? :)

OpenStudy (johnweldon1993):

Sure....think of the speed as 'x' The trip took the same amount of time...so they are equal to each other The current add 6mph the first half...and takes away 6mph the return half so Time * speed = distance right? so 4(x + 6) = 6(x - 6) 4x + 24 = 6x - 36 2x = 60 x = ?

OpenStudy (anonymous):

x=30

OpenStudy (queelius):

So, distance is equal to velocity * time. We know that he travelled for 4 hours in one direction, and 6 hours in the other, and the distance was the same in both directions. It took longer to go back, so when going back, he went against the current. So, 4(x+6) = 6(x - 6). Solve for x, where x is the velocity he would be travelling at in still water.

OpenStudy (johnweldon1993):

Right...x = 30 :)